# CRAFT-LoRA: Content-Style Personalization via Rank-Constrained Adaptation and Training-Free Fusion

### v1

@@ 1 Introduction
Recent advancements in deep learning have fueled remarkable progress in generative models, particularly within the domain of personalized image generation [11, 2, 30]. Unlike generic image synthesis, personalized generation focuses on customizing outputs based on user-provided references, such as a specific subject’s appearance or stylistic attributes. In more sophisticated scenarios, these systems aim to simultaneously preserve subject identity while emulating a target style, enabling applications in creative design, entertainment, digital avatars, personalized marketing, and beyond.
@@
Among the various techniques developed for personalized image generation, Low-Rank Adaptation (LoRA) has emerged as a particularly prominent and efficient approach [15]. LoRA facilitates adaptation by introducing small, trainable low-rank matrices into specific layers of large pre-trained text-to-image diffusion models, thereby allowing for efficient fine-tuning on new concepts using only a small number of references. When integrated with personalization frameworks like DreamBooth, LoRA-based methods enable the customization of generated images to specific content, e.g. preserving the identity of a person, or style, e.g. emulating a particular artistic rendering or texture, making them widely adopted [29]. Moreover, LoRA exhibits a unique compositional property: multiple independently trained LoRA modules, each corresponding to different visual concepts, can be simultaneously applied during inference. This enables the model to generate images that blend diverse attributes, such as combining a specific subject with a distinct artistic style. However, the combinatorial use of LoRA weights introduces new challenges. Naively merging multiple LoRA modules often leads to degraded visual quality or semantic inconsistency, due to entangled or conflicting latent representations. Consequently, recent works have explored various strategies for more effective LoRA fusion, including approaches that re-optimize weights for specific combinations [31], adjust internal attention mechanisms [9], or dynamically modulate different LoRAs’ influence during generation [22].
@@
Despite recent progress, existing LoRA combination strategies still face several fundamental challenges, which motivate the improvements proposed in this paper. First, although pre-trained diffusion models empirically exhibit some degree of composability with respect to LoRA modules, these models are not explicitly trained to support such combinations. As a result, directly merging LoRA weights often fails to achieve a clean disentanglement between content and style factors present in the reference images [19, 18, 1]. Inspired by the meta-learning paradigm, particularly Model-Agnostic Meta-Learning (MAML) [8, 6], which seeks optimal initializations for fast and generalizable task adaptation, we propose a novel network re-initialization strategy for pre-trained diffusion models that enhances their compatibility with compositional LoRA usage. Specifically, we introduce a low-rank constrained adaptation phase over the parameters of the pre-trained diffusion model inspired by PaRa [5]. The additional training stage narrows the model’s generative space, shifting it from a broadly generalized prior to a more compact and disentangled representation space that better supports the fusion of multiple LoRA modules.
Another critical limitation lies in the inherent granularity ambiguity associated with the definition of a “subject” in personalized image generation. In practical scenarios, the notion of a subject often encompasses a hierarchy of visual attributes—for instance, a portrait image may include multiple levels of subject-specific features, e.g. facial identity, hairstyle, clothing style, and accessory details. However, current methods typically collapse this rich semantic structure into a single coarse-grained token representation within the input prompt [24, 20]. This oversimplification neglects the internal structure of subject attributes and lacks the mechanisms to differentiate or control the preservation and variation of fine-grained elements.
@@
To address this, we propose a novel expert encoder system that enables fine-grained control over subject features at different levels. Our system implements specially designed content and style adapters operating in disjoint network layers. The expert encoder processes identity labels, content descriptions, and style specifications through separate branches, facilitating selective weights aggregation based on semantic relevance The designed expert system enables precise delineation of feature preservation hierarchies through prompts, allowing for discriminative activation of specific subject attributes at variable granularity levels according to specified semantic constraints.
@@
Finally, existing LoRA fusion strategies, such as ZipLoRA [31], typically rely on additional optimization procedures to reconcile differences between LoRA modules [37, 12]. While effective to some extent, directly modifying LoRA parameters during fusion can inadvertently alter or suppress critical elements encoded in the original weights, leading to loss of identity or stylistic fidelity. Moreover, these methods incur non-trivial computational overhead. To overcome these limitations, we propose a novel tuning-free LoRA combination method that circumvents the need for parameter-level fusion. At the core of our approach is our developed asymmetric Classifier-Free Guidance (CFG) mechanism applied during the diffusion process. Rather than adjusting LoRA weights directly, we selectively apply time-dependent modifications only to the conditional path where combined LoRA weights are applied, while keeping the unconditional path that uses only the base model without LoRA adapters fixed to the base model weights. Our asymmetric design yields a dynamic guidance signal that effectively steers the generation process toward a desirable balance between content and style, without introducing additional training cost or risking the degradation of essential features, thereby offering a lightweight and robust alternative to existing LoRA fusion techniques.
@@
As illustrated in Figure 2, our proposed unified framework integrates rank-constrained fine-tuning to learn decoupled content and style subspaces, leverages prompt-guided mechanisms for semantic control and extension via selective adapter aggregation guided by an expert encoder, and employs a time-step-aware guidance correction scheme for stable, high-fidelity generation without extra training.
@@ 2.1 Personalized image generation
2.2 Stylization through Fine-Tuning Methods
The rapid advancement of text-to-image diffusion models has spurred diverse personalization strategies. Full-model fine-tuning approaches, exemplified by DreamBooth [27], can achieve high-fidelity style transfer but incur substantial computational costs and require lengthy retraining on domain-specific examples. To mitigate these demands, Parameter-Efficient Fine-Tuning techniques have emerged [28]. Among these, LoRA introduces small and trainable low-rank decomposition matrices, injecting new stylistic or semantic concepts while keeping the majority of the pre-trained model weights frozen, thus striking a balance between adaptability and efficiency. Complementing weight-space adaptation methods, token-based optimization techniques, such as Textual Inversion [10], learn new embeddings to represent novel concepts within the model’s token vocabulary. Adapter-based schemes offer an alternative pathway, for example, StyleDrop [33] appends lightweight layers to attention blocks, isolating style information in dedicated modules. Meta-learning, as a cross-task method for domain transfer, is similar in principle to style transfer in personalized image generation, providing new inspirations for fine-tuning methods [39].
2.3 Content-Style LoRAs Merging
Methods for combining multiple LoRA modules, effective for capturing distinct concepts, are explored to achieve simultaneous content and style adaptation. Parameter interpolation is a basic approach, which is prone to suboptimal outcomes, especially when parameters conflict [34, 36]. To address these challenges, several studies focus on refining the composition mechanism or the application process itself. ZipLoRA [31] learns column-wise mixing coefficients to better balance style and content elements, although it introduces the need for additional optimization procedures for each new content-style pairing. B-LoRA [9], in contrast, differentiates the functional roles within attention modules, aiming for a more effective separation of object and style representations. LoRA Composition [40] employs a cyclic update mechanism, allowing LoRA weights to iteratively guide the generation without explicit parameter merging.
@@ 3.1 Preliminaries
LoRA is an efficient fine-tuning method that adapts pretrained diffusion models by adding trainable low-rank matrices $A\in\mathbb{R}^{r\times n}$ and $B\in\mathbb{R}^{m\times r}$ to frozen weights $W_{0}\in\mathbb{R}^{m\times n}$, yielding $W=W_{0}+BA$ with $r\ll\min(m,n)$. In personalized image generation, content LoRAs $\Delta W_{c}$ represent subject-specific features like identity and structure, while style LoRAs $\Delta W_{s}$ capture artistic aspects such as texture and color. Both can be trained efficiently from a single reference image, and during inference they can be composed into the backbone to form personalized generative elements.
@@ 3.2 Framework Overview
We propose CRAFT-LoRA, a unified framework for personalized image synthesis that addresses the challenges of content-style disentanglement and fusion in LoRA-based methods. As shown in Fig. 2, the framework includes three main stages: (1) Training Stage, where content and style LoRA weights are learned separately based on a rank-constrained fine-tuned backbone; (2) Prompt Text Guidance, which employs an expert encoder system with specialized branches to provide semantic control over content and style during training and inference; (3) Inference Stage, utilizing a timestep-aware asymmetric classifier-free guidance scheme for stable and flexible fusion of LoRA weights without extra training.
@@ 3.3 Rank-Limited Fine-Tuning of Backbone
As mentioned in the introduction, pre-trained diffusion models do not explicitly support clear disentanglement of content and style factors when adapting weights from a reference image, thus separation enhancement requires additional fine-tuning of the diffusion backbone. Motivated by the adaptability of MAML [8, 17] under few-shot constraints and leveraging the rich hierarchical features of SDXL [25], we propose to fine-tune the U-Net backbone under a rank-constrained mechanism. Our approach provides optimized initialization and structurally improves the disentanglement of content and style in the U-Net architecture, providing a more effective foundation for subsequent LoRA separation and personalized synthesis.
3.3.1 Rank Reduction Fine-Tuning.
Inspired by PaRa method [5], we denote the frozen backbone weight matrix of layer $l$ as $W_{l}^{(0)}\in\mathbb{R}^{d_{\mathrm{out}}^{l}\times d_{\mathrm{in}}^{l}}$, with $L$ layers in total. We introduce a learnable basis matrix $B_{l}\in\mathbb{R}^{d_{\mathrm{in}}^{l}\times r_{l}}$ per layer, where $r_{l}$ is the rank constraint satisfying $r_{l}\ll\min(d_{\mathrm{out}}^{l},d_{\mathrm{in}}^{l})$. Computing the QR decomposition $B_{l}=Q_{l}R_{l}$ yields an orthonormal basis $Q_{l}$ with $Q_{l}^{\top}Q_{l}=I$. The rank-limited fine-tuning projects out components of $W_{l}^{(0)}$ along $Q_{l}$, yielding updated weights:
@@
This constrains updates to directions orthogonal to the learned low-rank subspace, introducing an architectural bias that encourages reduced overlap between content and style representations. We show empirically in Appendix. Appendix that it reduces cross-influence.
@@
assigning higher ranks to earlier layers, which typically encode low-level structural and textural information where content and style factors are more intertwined, thus requiring more adaptation capacity.
For content-style disentanglement, we train separate basis matrices $B_{l}^{\mathrm{content}}$ and $B_{l}^{\mathrm{style}}$, obtaining corresponding orthonormal subspaces $Q_{l}^{\mathrm{content}}$ and $Q_{l}^{\mathrm{style}}$. Merging these via concatenation and QR decomposition and compute the combined weight update:
@@
3.3.2 Contrastive Content-Style Pairs.
To facilitate effective disentanglement of content and style during fine-tuning, we construct a carefully curated dataset consisting of $\mathbf{100}$ contrastive content-style image pairs. These pairs are specifically designed so that either the content or the style remains constant while the other varies, explicitly guiding the model to differentiate and separately capture content and style characteristics.
We operationalize the notion of content and style using frequency-domain decomposition: content corresponds primarily to low-frequency components encoding structural and semantic information such as object layouts, while style corresponds to high-frequency components capturing textures and color palettes. Specifically, we apply Gaussian low-pass filtering with cutoff frequency $\sigma=0.35$ (normalized) to extract content features, and compute the residual as style features. These frequency-separated representations are then used as conditioning signals during text-to-image generation to produce the contrastive pairs.
The content and style samples under this frequency method are shown in Figure 3.
Concretely, we select $\mathbf{10}$ distinct content reference images $\{C_{i}\}_{i=1}^{10}$ and $\mathbf{10}$ distinct style reference images $\{S_{j}\}_{j=1}^{10}$. Exhaustively pairing them yields:
@@
Fixing a content image $C_{1}$ and combining it with styles $\{S_{1},\dots,S_{10}\}$ produces 10 pairs that vary only in style, while fixing a style and varying content produces complementary pairs. These contrastive pairs are directly used to drive the rank-limited backbone training described above. For each pair, the two images are fed as separate inputs: one provides supervision for learning the content-specific basis $B_{l}^{\mathrm{content}}$, while the other supervises the style-specific basis $B_{l}^{\mathrm{style}}$. The corresponding orthonormal subspaces $Q_{l}^{\mathrm{content}}$ and $Q_{l}^{\mathrm{style}}$ are then merged, and the backbone is updated according to Eqs.3,4. In this way, the model allocates content and style variations to disjoint low-rank subspaces, ensuring that the backbone initialization $W_{l}^{\mathrm{combined}}$ encodes reduced interference between the two factors.
Empirically, combining rank-limited updates with contrastive pairs leads to a 12% reduction in average CLIP-based content–style similarity, indicating that the learned features exhibit reduced cross-influence. As further demonstrated in Table 6(b), the removal of rank-limited backbone fine-tuning results in clear drops in both content and style fidelity, quantitatively confirming its role in disentanglement. Details on dataset construction and training settings are provided in Appendix Appendix.
Figure 3: Separating content and style components in a single image through frequency domain decomposition
@@ 3.4 Specialized Decoupling and Fusion
Building on the rank-limited backbone initialization, we introduce a prompt-guided framework to further separate content and style at the adapter level. During training, content and style LoRA updates are learned on disjoint layer subsets with explicit prompt markers. At inference, an expert encoder routes prompt information to selectively activate these adapters, enabling flexible and fine-grained fusion.
3.4.1 Prompt-Guided Decoupling and Encoding.
Given a raw prompt $p$ containing explicit markers for content and style, e.g. p=p=”A photo of a person <c>smiling in a watercolor style <s>”, we treat these markers as routing signals to regulate LoRA module activation and to prevent leakage of content semantics into the style branch and vice versa. The prompt with markers removed is encoded by the SDXL text encoder to produce a unified semantic embedding:
@@
Following the standard LoRA parametrization, each layer $i$ is updated by low-rank matrices. To bias the network towards disentanglement, we allocate disjoint index sets $I_{c}$ and $I_{s}$ corresponding to content- and style-sensitive layers: lower and middle layers primarily capture structure and identity (content), while higher layers encode textures and rendering patterns (style). Content-specific matrices $A_{i}^{(c)}$ and style-specific matrices $A_{i}^{(s)}$ are trained with the semantic embedding, and updates are restricted as
@@
where $W_{\text{init}}$ denotes the rank-limited backbone obtained in Eqs.3,4. The operator $\mathsf{E}_{i}(\cdot)$ injects its input matrix into the slot of the $i$-th layer while filling zeros elsewhere, so the summation is a layerwise direct sum rather than a cross-layer addition.
If markers <c> or <s> are absent, the corresponding branch remains inactive. Users may also directly adjust $\gamma_{c},\gamma_{s}$ to control the relative influence of content and style. This enables flexible compositions such as “preserve content but replace style,” without retraining. The full inference procedure is summarized in Algorithm 1, showing how prompts route into disjoint LoRA branches and produce the aggregated backbone $W_{\mathrm{agg}}$.
@@
3.4.2 Asymmetric Classifier-Free Guidance
In standard classifier-free guidance (CFG), the conditional and unconditional paths share the same model weights. When LoRA updates are applied, this causes the unconditional path to be contaminated by content and style adapters, leading to instability in image generation. To address this issue, we propose Asymmetric Classifier-Free Guidance (ACFG), where the conditional path uses LoRA-adapted weights while the unconditional path remains anchored to the rank-limited backbone $W_{\text{init}}$. This asymmetry yields a cleaner and more stable guidance signal without requiring additional training.
Let $i\in\{1,\ldots,L\}$ index model layers and $t\in\{1,\ldots,T\}$ diffusion timesteps. Instead of uniformly applying LoRA updates at all timesteps, we introduce time-dependent activation schedules for content and style: $\gamma_{c}(t)=\mathbb{1}_{t\in T_{c}},\gamma_{s}(t)=\mathbb{1}_{t\in T_{s}}.$ The conditional weights at timestep $t$ are defined as
where $W^{\text{init}}_{i}$ denotes the rank-limited backbone, and $\Delta W_{i}^{(c)},\Delta W_{i}^{(s)}$ are content and style LoRA updates. While the unconditional path uses only the initialization backbone:
$W_{i}^{\mathrm{uncond}}(t)=W^{\text{init}}_{i},\quad\forall\,t,i.$ (10)
The conditional prediction is $\epsilon_{\mathrm{cond}}=\epsilon_{\theta}(x_{t}\mid W^{\mathrm{cond}}(t),\mathrm{cond})$, and the unconditional prediction is $\epsilon_{\mathrm{uncond}}=\epsilon_{\theta}(x_{t}\mid W^{\mathrm{uncond}},\varnothing)$. Then the final guided estimate is
where $\omega$ controls guidance strength. By keeping the unconditional path free of LoRA updates, ACFG avoids interference between content and style adapters and provides a more reliable guidance signal.
Figure 5: Extended visual results. (a) Content–style generations augmented with additional prompt descriptions e.g. “catching a frisbee”, “wearing a hat”, and “driving a car”. The results show that our method flexibly integrates dynamic semantics while preserving the specified content and style. (b) Single-branch generations where only the content LoRA or only the style LoRA is activated. Content-only generations preserve object identity across neutral renderings, while style-only generations transfer artistic features across different subjects, demonstrating disentanglement and controllability.

### v2

@@ 1 Introduction
Recent advancements in deep learning have fueled remarkable progress in generative models, particularly within personalized image generation [11, 2, 33]. Unlike generic image synthesis, personalized generation focuses on customizing outputs based on user-provided references, such as a specific subject’s appearance or stylistic attributes, enabling applications in creative design, entertainment, digital avatars, and personalized marketing.
@@
Among the various techniques developed for personalized image generation, Low-Rank Adaptation (LoRA) [15] has emerged as a particularly efficient approach by introducing small, trainable low-rank matrices into pre-trained text-to-image diffusion models, allowing fine-tuning on new concepts using only a small number of references. When integrated with frameworks like DreamBooth, LoRA-based methods enable customization to specific content or style from minimal examples [32]. Moreover, multiple independently trained LoRA modules can be simultaneously applied during inference, enabling images that blend diverse attributes such as combining a specific subject with a distinct artistic style. However, naively merging multiple LoRA modules often leads to degraded quality or semantic inconsistency due to entangled representations, motivating recent work on more effective fusion strategies [34, 9, 25].
@@
Despite recent progress, existing LoRA combination strategies still face several fundamental challenges. First, although pre-trained diffusion models empirically exhibit some degree of composability with respect to LoRA modules, these models are not explicitly trained to support such combinations. As a result, directly merging LoRA weights often fails to achieve clean disentanglement between content and style factors [21, 20, 1]. Inspired by the meta-learning paradigm, particularly MAML [8, 6], which seeks optimal initializations for fast and generalizable task adaptation, we introduce a low-rank constrained adaptation phase inspired by PaRa [5] that narrows the model’s generative space to a more compact and disentangled representation that better supports multi-LoRA fusion.
@@
Second, the notion of a subject in personalized image generation often encompasses a hierarchy of visual attributes, e.g. facial identity, hairstyle, clothing style, and accessory details. However, current methods typically collapse this rich semantic structure into a single coarse-grained token representation [27, 22], neglecting the internal structure of subject attributes and lacking mechanisms to control the preservation of fine-grained elements. To address this, we propose an expert encoder system with specially designed content and style adapters operating in disjoint network layers. The expert encoder processes identity labels, content descriptions, and style specifications through separate branches, facilitating selective aggregation based on semantic relevance and enabling fine-grained control over feature preservation at different levels.
@@
Third, existing LoRA fusion strategies such as ZipLoRA [34] typically rely on additional optimization procedures to reconcile differences between LoRA modules [41, 12]. Directly modifying LoRA parameters during fusion can inadvertently alter or suppress critical elements encoded in the original weights, leading to loss of identity or stylistic fidelity, while also incurring non-trivial computational overhead. To overcome these limitations, we propose a tuning-free asymmetric Classifier-Free Guidance (CFG) mechanism. Rather than adjusting LoRA weights directly, we selectively apply time-dependent modifications only to the conditional path while keeping the unconditional path anchored to base model weights, yielding a dynamic guidance signal that steers generation toward a desirable content-style balance without additional training cost.
@@
As illustrated in Figure 2, our proposed unified framework integrates rank-constrained fine-tuning to learn decoupled content and style subspaces, leverages prompt-guided mechanisms for semantic control via selective adapter aggregation guided by an expert encoder, and employs a timestep-aware guidance correction scheme for stable, high-fidelity generation without extra training.
@@ 2.1 Personalized image generation
2.2 Fine-Tuning and LoRA Composition for Stylization
The rapid advancement of text-to-image diffusion models has spurred diverse personalization strategies. Full-model fine-tuning approaches such as DreamBooth [30] achieve high-fidelity transfer but incur substantial computational costs and require lengthy retraining on domain-specific examples. To mitigate these demands, Parameter-Efficient Fine-Tuning (PEFT) techniques have emerged [31, 12]. Among these, LoRA [15] introduces small, trainable low-rank decomposition matrices, injecting new stylistic or semantic concepts while keeping the majority of pre-trained weights frozen, thus striking a balance between adaptability and efficiency. Complementary approaches include token-based optimization techniques such as Textual Inversion [10], which learns new embeddings to represent novel concepts within the model’s token vocabulary, and adapter-based schemes such as StyleDrop [37], which appends lightweight layers to attention blocks to isolate style information in dedicated modules. Meta-learning, as a cross-task method for domain transfer, shares similar principles with style transfer in personalized image generation, providing new inspirations for fine-tuning methods [45].
A key challenge arises when combining multiple LoRA modules for simultaneous content and style adaptation. Parameter interpolation [38, 40] provides a basic approach but is prone to suboptimal outcomes, especially when parameters conflict. To address this, several studies refine the composition mechanism: ZipLoRA [34] learns column-wise mixing coefficients for each content-style pairing, B-LoRA [9] differentiates functional roles within attention modules for content-style separation, and LoRA Composition [46] employs a cyclic update mechanism to iteratively guide generation without explicit parameter merging. Training-free approaches have also been explored, with K-LoRA [25] selectively applying LoRAs to different attention layers and CLoRA [24] proposing contrastive test-time composition. Other directions include QR-LoRA [43], which applies QR decomposition for parameter-efficient training, and LoRA.rar [35], which learns merging via hypernetworks. While these methods improve the merging or composition process, the upstream entanglement inherent in standard backbones remains underexplored.
@@ 3.1 Preliminaries
LoRA is an efficient fine-tuning method that adapts pre-trained diffusion models by adding trainable low-rank matrices $A\in\mathbb{R}^{r\times n}$ and $B\in\mathbb{R}^{m\times r}$ to frozen weights $W_{0}\in\mathbb{R}^{m\times n}$, yielding $W=W_{0}+BA$ with $r\ll\min(m,n)$. In personalized image generation, content LoRAs $\Delta W_{c}$ represent subject-specific features such as identity and structure, while style LoRAs $\Delta W_{s}$ capture artistic aspects such as texture and color. Both can be trained efficiently from a single reference image, and during inference they can be composed into the backbone to produce personalized outputs.
@@ 3.2 Framework Overview
We propose CRAFT-LoRA, a unified framework for personalized image synthesis that addresses content-style disentanglement and fusion in LoRA-based methods. As shown in Fig. 2, the framework consists of three main stages: (1) Training Stage, where content and style LoRA weights are learned separately based on a rank-constrained fine-tuned backbone; (2) Prompt Text Guidance, which employs an expert encoder system with specialized branches to provide semantic control over content and style during training and inference; (3) Inference Stage, utilizing a timestep-aware asymmetric classifier-free guidance scheme for stable and flexible fusion of LoRA weights without extra training.
@@ 3.3 Rank-Limited Fine-Tuning of Backbone
As discussed in the introduction, pre-trained diffusion models do not explicitly support clear disentanglement of content and style factors when adapting weights from a reference image. Motivated by the adaptability of MAML [8, 18] under few-shot constraints and leveraging the rich hierarchical features of SDXL [28], we propose to fine-tune the U-Net backbone under a rank-constrained mechanism that provides optimized initialization and structurally improves content-style disentanglement for subsequent LoRA training.
Rank Reduction Fine-Tuning. Inspired by PaRa [5], we denote the frozen backbone weight matrix of layer $l$ as $W_{l}^{(0)}\in\mathbb{R}^{d_{\mathrm{out}}^{l}\times d_{\mathrm{in}}^{l}}$, with $L$ layers in total. We introduce a learnable basis matrix $B_{l}\in\mathbb{R}^{d_{\mathrm{in}}^{l}\times r_{l}}$ per layer, where $r_{l}$ is the rank constraint satisfying $r_{l}\ll\min(d_{\mathrm{out}}^{l},d_{\mathrm{in}}^{l})$. Computing the QR decomposition $B_{l}=Q_{l}R_{l}$ yields an orthonormal basis $Q_{l}$ with $Q_{l}^{\top}Q_{l}=I$. The rank-limited fine-tuning projects out components of $W_{l}^{(0)}$ along $Q_{l}$, yielding updated weights:
@@
This operation constrains updates to directions orthogonal to the learned low-rank subspace, introducing an architectural bias that encourages reduced overlap between content and style representations. We show empirically in Appendix Appendix that it reduces cross-influence.
@@
assigning higher ranks to earlier layers, which typically encode low-level structural and textural information where content and style factors are more intertwined, thus requiring more adaptation capacity. In implementation, we set $r_{\max}=128$ and $r_{\min}=4$. Compared to a uniform rank $r=64$ across all layers, this schedule improves Content Similarity by 4.8%, Style Similarity by 5.2%, and reduces cross-influence by 11.3%, confirming the benefit of hierarchical rank allocation.
For content-style disentanglement, we train separate basis matrices $B_{l}^{\mathrm{content}}$ and $B_{l}^{\mathrm{style}}$, obtaining corresponding orthonormal subspaces $Q_{l}^{\mathrm{content}}$ and $Q_{l}^{\mathrm{style}}$. Merging these via concatenation and QR decomposition yields the combined weight update. Specifically, we first concatenate the basis matrices $B_{l}^{\mathrm{content}}\mid B_{l}^{\mathrm{style}}]$ and apply QR decomposition to their corresponding orthonormal subspaces:
@@
Figure 3: Content-style separation via frequency domain decomposition. Each group shows an original image with its frequency-separated elements and additional samples from the same group. Left: content group, where low-frequency components capture structural and semantic information. Right: style group, where high-frequency components capture textures and artistic rendering.
Contrastive Content-Style Pairs. To facilitate effective disentanglement during fine-tuning, we construct a curated dataset of $\mathbf{100}$ contrastive content-style image pairs. These pairs are designed so that either the content or the style remains constant while the other varies, guiding the model to separately capture content and style characteristics.
We operationalize the notion of content and style using frequency-domain decomposition: content corresponds primarily to low-frequency components encoding structural and semantic information such as object layouts, while style corresponds to high-frequency components capturing textures and color palettes. Specifically, we apply Gaussian low-pass filtering with cutoff frequency $\sigma=0.35$ (normalized) to extract content features, and compute the residual as style features. The cutoff was selected empirically by evaluating separation quality across $\sigma\in\{0.2,0.25,0.3,0.35,0.4,0.45,0.5\}$ via MLLM-based assessment, with $\sigma=0.35$ yielding the best content-style separation across diverse reference images.
The content and style samples under this frequency method are shown in Figure 3. Concretely, we select $\mathbf{10}$ distinct content reference images $\{C_{i}\}_{i=1}^{10}$ and $\mathbf{10}$ distinct style reference images $\{S_{j}\}_{j=1}^{10}$. Exhaustively pairing them yields:
@@
Fixing a content image $C_{1}$ and combining it with styles $\{S_{1},\dots,S_{10}\}$ produces 10 pairs that vary only in style, while fixing a style and varying content produces complementary pairs.
These contrastive pairs are directly used to drive the rank-limited backbone training described above. For each pair, the two images are fed as separate inputs: one provides supervision for learning the content-specific basis $B_{l}^{\mathrm{content}}$, while the other supervises the style-specific basis $B_{l}^{\mathrm{style}}$. The corresponding orthonormal subspaces $Q_{l}^{\mathrm{content}}$ and $Q_{l}^{\mathrm{style}}$ are then merged, and the backbone is updated according to Eqs.3,4. In this way, the model allocates content and style variations to disjoint low-rank subspaces, ensuring that the backbone initialization $W_{l}^{\mathrm{combined}}$ encodes reduced interference between the two factors. We find that diversity matters more than quantity since 50 pairs already achieve 95% of full performance, while increasing to 200 pairs yields less than 2% additional improvement.
Empirically, combining rank-limited updates with contrastive pairs leads to a 12% reduction in average CLIP-based content–style similarity, indicating that the learned features exhibit reduced cross-influence. As further demonstrated in Table 6(b), removing Rank-FT results in the largest performance drop among all components, quantitatively confirming its central role in disentanglement. Details on dataset construction and training settings are provided in Appendix Appendix.
@@ 3.4 Specialized Decoupling and Fusion
Building on the rank-limited backbone initialization, we introduce a prompt-guided framework to further separate content and style at the adapter level. During training, content and style LoRA updates are learned on disjoint layer subsets with explicit prompt markers. At inference, an expert encoder routes prompt information to selectively activate these adapters, enabling flexible fusion and fine-grained control.
Prompt-Guided Decoupling and Encoding. Given a raw prompt $p$ containing explicit markers for content and style, e.g. $p=$ “A photo of a person <c> smiling in a watercolor style <s>”, we treat these markers as routing signals to regulate LoRA module activation and to prevent leakage of content semantics into the style branch and vice versa. The prompt with markers removed is encoded by the SDXL text encoder to produce a unified semantic embedding:
@@
Following standard LoRA parametrization, each layer $i$ is updated by low-rank matrices. To bias the network towards disentanglement, we allocate disjoint index sets $I_{c}$ and $I_{s}$: lower and middle layers primarily capture structure and identity (content), while higher layers encode textures and rendering patterns (style). Content-specific matrices $A_{i}^{(c)}$ and style-specific matrices $A_{i}^{(s)}$ are trained with the semantic embedding, and updates are restricted as
@@
where $W_{\text{init}}$ denotes the rank-limited backbone obtained in Eqs.3,4. The operator $\mathsf{E}_{i}(\cdot)$ injects its input into the $i$-th layer while filling zeros elsewhere, so the summation operates as a layerwise direct sum.
If markers <c> or <s> are absent, the corresponding branch remains inactive. Users may also directly adjust $\gamma_{c},\gamma_{s}$ to control the relative influence of content and style, enabling flexible compositions such as “preserve content but replace style” without retraining. The full inference procedure is summarized in Algorithm 1, showing how prompts route into disjoint LoRA branches and produce the aggregated backbone $W_{\mathrm{agg}}$.
Asymmetric Classifier-Free Guidance In standard classifier-free guidance (CFG), the conditional and unconditional paths share the same model weights. When LoRA updates are applied, this causes the unconditional path to be contaminated by content and style adapters, leading to instability in image generation. To address this issue, we propose Asymmetric Classifier-Free Guidance (ACFG), where the conditional path uses LoRA-adapted weights while the unconditional path remains anchored to the rank-limited backbone $W_{\text{init}}$.
Let $i\in\{1,\ldots,L\}$ index model layers and $t\in\{1,\ldots,T\}$ diffusion timesteps. Instead of uniformly applying LoRA updates at all timesteps, we introduce time-dependent activation schedules $\gamma_{c}(t)=\mathbb{1}_{t\in T_{c}}$ and $\gamma_{s}(t)=\mathbb{1}_{t\in T_{s}}$, where $T_{c}$ and $T_{s}$ denote the active timestep ranges for content and style respectively. In practice, we activate content LoRA during early-to-mid timesteps to establish structural layout, and style LoRA during mid-to-late timesteps to refine textures and rendering, reflecting the coarse-to-fine nature of the diffusion process. The conditional weights at timestep $t$ are:
while the unconditional path uses only the initialization backbone throughout:
$W_{i}^{\mathrm{uncond}}(t)=W^{\text{init}}_{i},\quad\forall\,t,i.$ (10)
The conditional and unconditional noise predictions are $\epsilon_{\mathrm{cond}}=\epsilon_{\theta}(x_{t}\mid W^{\mathrm{cond}}(t),\mathrm{cond})$ and $\epsilon_{\mathrm{uncond}}=\epsilon_{\theta}(x_{t}\mid W^{\mathrm{uncond}},\varnothing)$ respectively. The final guided estimate is:
where $\omega$ controls guidance strength. The key insight is that by anchoring the unconditional path to $W_{\text{init}}$, the guidance signal $\epsilon_{\mathrm{cond}}-\epsilon_{\mathrm{uncond}}$ isolates the effect of LoRA adapters at each timestep, rather than conflating adapter influence with the unconditional baseline. Combined with the time-dependent schedules, this allows content structure to be established before style details are applied, reducing interference between the two factors. Crucially, ACFG requires no additional training or parameter optimization and maintains the same two-pass computation as standard CFG. Therefore, ACFG can be directly applied to standard SDXL LoRAs as a single module without the rank-limited backbone, providing improved fusion stability albeit with reduced disentanglement.
@@
Figure 5: Extended visual results. (a) Content–style generations augmented with additional prompt descriptions e.g. “catching a frisbee”, “wearing a hat”, and “driving a car”. Our method flexibly integrates dynamic semantics while preserving the specified content and style. (b) Single-branch generations where only the content LoRA or only the style LoRA is activated. Content-only generations preserve object identity across neutral renderings, while style-only generations transfer artistic features across different subjects.